Evaluate
h(x) = x2 – 4x + 5
h(4) =
h(–4) =
h(0) =

Found 2 solutions by solver91311, JBarnum:
Answer by solver91311(24713)



If



Then



So




and



Congratulations, you get to do your very own arithmetic.

John


Answer by JBarnum(2146)

simply replace the x with the value given and solve
h(x)= x^2–4x+5
h(4)= (4^2)–4(4)+5 = 16-16+5 = 5
h(–4)= (-4)^2–4(-4)+5 = 16+16+5 = 37
h(0)= 0^2–4(0)+5 = 5